OpenStudy (anonymous):

One solution of the equation 3x2 - 16x + 5 = 0 is 1/3 . What is the other solution?

OpenStudy (anonymous):

(3x-1)(x-5) = 0 3x-1 = 0 => x = 1/3 x-5 = 0 => x = 5